Please help me solve this story problem........ if the measure of one angle of a triangle is 3 times another and the third angle is twenty degrees larger than the measure of the smaller of the other two angles what is the measure of all the angles of the triangle
Answer by jorel1380(3719)

n+3n+(n+3n+20)=180
8n+20=180
8n=160
n=20
3n=60
4n+20=100
The angles are 20 degrees, 60 degrees, and 100 degrees, respectively..